# Hanging on the cliff: EMRI formation with local two-body relaxation and post-Newtonian dynamics

### v1

@@ 4.1 EMRI-to-DP ratio and cliffhanger EMRIs
Consider now $a_{\alpha}$ and $a_{\beta}$ as the values of the semi-major axis of the orbit at two subsequent generic apocentres. For each run that resulted in the formation of an EMRI, we consider the largest ratio $a_{\alpha}/a_{\beta}$ to set a criterion to distinguish cliffhanger and traditional EMRIs. Consistently with the reasoning that led us to Eq. (35), we determine the passage across the cliffhanger region if at some point in the orbital evolution the semi-major axis is cut in half or more during a single period, thus for all the runs where $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}>2$. This threshold is somehow arbitrary, but offers a quantitative way to identify cliffhanger EMRIs. Figure 10 shows the result of this investigation for $M_{\bullet}=10^{5}\,\mathrm{M_{\sun}}$ and $M_{\bullet}=10^{6}\,\mathrm{M_{\sun}}$. For both MBH masses, all EMRIs forming for $a_{0}\ll 10^{-2}R_{\mathrm{inf}}$ are “classical” EMRIs. Both cases show an increase in the average value of $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}$ for larger $a_{0}$, but only for $M_{\bullet}=10^{5}\,\mathrm{M_{\sun}}$ the threshold of 2 is crossed. Moreover, we verify that the prediction for the largest possible $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}$ ratio (which occurs at the loss cone edge) given in Eq. (38) is verified for $M_{\bullet}=10^{5}\,\mathrm{M_{\sun}}$ for large $a_{0}$, with some deviations for smaller $a_{0}$. This is likely because the approximation given in Eq. (37) is less precise for smaller semi-major axes.
@@ 4.2 Local versus orbit-averaged approach
The relaxation timescale for a particle with energy $\mathcal{E}$ and pericentre $r_{-}$ estimates the average time that the particle passes on the orbit before being scattered away. When $r_{-}$ is much smaller than the semi-major axis of the orbit, it estimates the time before $r_{-}$ is changed by the order of itself, as energy remains almost constant (Cohn & Kulsrud 1978). In this small $r_{-}$ regime the relaxation timescale for the orbit can be estimated as $t_{\mathrm{rlx}}\simeq P\,r_{-}/(q\,r_{\mathrm{lc}})$ (Merritt 2013a; Bortolas et al. 2023; Broggi et al. 2024), where $P$ is the orbital period. Consider now an orbit with energy $\mathcal{E}$ and $r_{-}<r_{\mathrm{lc}}$: in a time $P$ it will be reached by a number $\mathrm{d}N$ of stellar objects due to two-body relaxation. In the orbit-averaged approach, they will all be DPs produced in the subsequent orbital period, with a rate from that orbit of $\mathrm{d}\dot{N}=\mathrm{d}N/P$. In the local approach, if $t_{\mathrm{rlx}}<P$, only the stellar objects that will take less than $t_{\mathrm{rlx}}$ to reach the pericentre will be DPs, and the corresponding rate will be $\mathrm{d}\dot{N}\simeq\mathrm{d}N\,t_{\mathrm{rlx}}/P^{2}=r_{-}/(q\,r_{\mathrm{lc}})\,\mathrm{d}N/P$. Therefore, the orbit-averaged approach will overestimate the rate of DPs for $q\gtrsim 1$ due to the “frozen orbit” approach, as they have $r_{-}<r_{\mathrm{lc}}$ by definition. The same reasoning applies to cliffhangers, that are produced at very small pericentres, too. However, being their pericentre larger than $r_{\mathrm{lc}}$ by definition, the overestimation of cliffhangers in the orbit-averaged assumption will be lower. Consequently, we expect that the orbit-averaged approach will underestimate $S(a)$.
@@
In order to test this expectation we re-ran the three sets of simulations with $M_{\bullet}=10^{4},10^{5},3\times 10^{5}\mathrm{M}_{\sun}$ with the exact same system parameters but employing the orbit-averaged procedure described in Sec. 3.3. Results of these test runs are compared to our default simulations in Fig. 12, and appear to confirm our expectation. The $S(a_{0})$ functions computed in the local vs orbit-averaged simulations nicely match up to the value of $a_{0}/R_{\rm inf}$ for which $q=1$, which marks the transition from empty to full loss cone (vertical dotted lines in figure). At larger $a_{0}$, the orbit-averaged approach tends to produce less cliffhanger EMRIs and more plunges, thus resulting in a lower $S(a_{0})$. Note that the trend becomes more severe by lowering the central MBH mass, consistent with the fact that the cliffhanger region becomes larger and larger. In practice, the smaller the central MBH, the bigger gets the cliffhanger region compared to the loss cone, and it becomes increasingly more difficult for the orbit-averaged approximation to capture the detailed dynamics of particle scattered within the cliffhanger region in real time along the orbit close to the pericentre, before getting to the loss cone, thus leading to an over-estimation of DPs.
@@
Figure 12: Comparison of the $S(a_{0})$ function obtained by employing the local (solid lines) vs orbit-averaged (dashed lines) treatment of two-body relaxation. Vertical dotted lines mark, for each case, the value of $a_{0}/R_{\rm inf}$ at which $q=1$, thus marking the transition from the empty (at smaller $a_{0}$) to the full loss cone regime. Crosses indicate points for which the orbit-averaged procedure explained in Sect. 3.3 failed many times.
4.3 Detailed investigation of the differences with QS23
@@
Our main results display some deviations with respect to other works in the literature. In particular the emergence of cliffhanger EMRIs and their abundance does not match findings by QS23. This is not surprising, given the many differences between our simulations and theirs. In particular, we simulate a two-population system, including the effect of solar-mass stars and stellar-mass BHs in the computation of the diffusion coefficient, we adopt a local treatment of two-body relaxation and we evolve the stellar-mass BH using PN equations of motion. Conversely, in QS23 scattering comes from a single population of solar-mass object, is treated as orbit-averaged and the dynamics of the stellar-mass BH is Newtonian.
@@
In the previous section, we already highlighted the importance of the local treatment of relaxation, which leads to a much larger relative fraction of EMRIs in the full loss cone regime. This can certainly account for some of the differences with QS23. For example in the case of $M_{\bullet}=10^{4}\mathrm{M}_{\sun}$, the orbit-averaged tests shown in Fig. 12 result in $S(a_{0})\approx 0.4$ at large $a_{0}$, which is much closer to the value of $\approx 0.25$ found in QS23 compared to the $\gtrsim 0.6$ found in our default simulations employing a local treatment of relaxation.
@@
In order to further test how the differences between the QS23 and our studies affect the results, we perform further targeted simulations, focusing on the $M_{\bullet}=3\times 10^{5}\mathrm{M}_{\sun}$ case, as it is close to the MBH mass value at which the majority of LISA EMRI detections is expected (Babak et al. 2017). We try to reproduce the results by QS23111111QS23 do not present results for a $M_{\bullet}=3\times 10^{5}\mathrm{M}_{\sun}$ scenario, however the authors gently agreed to share such results with us and let us show them in Fig. 13. by running our simulations with a setup which is as close as possible to theirs. We use a single population of scattering objects of $m_{\star}=1\mathrm{M}_{\sun}$, distributed according to a Denhen density profile characterized by $M_{\mathrm{tot},\star}=1425M_{\bullet}$, $r_{\mathrm{a},\star}=223.37R_{\mathrm{inf}}$ and $\gamma_{\star}=1.75$. With these parameters, our density profile departs from the single $r^{-1.75}$ power-low simulated by QS23 by less than 1% within the MBH influence radius.
@@
QS23 use Newtonian dynamics, and take into account emission of GWs through the equations detailed in Peters (1964). We do the same by setting $H_{1}=H_{2}=H_{2.5}=0$ and adding the contributions
@@
We report the results of such tests in Fig. 13.
Figure 13: Function $S(a_{0})$ given different assumptions on the dynamics and treatment of two-body relaxation. For the dynamics, we distinguish between Newtonian (N) or PN evolution of the system, and between employing the equations detailed in Peters (1964) or the 2.5PN term to treat loss of energy and angular momentum due to GWs emission. Instead, two-body relaxation can be accounted for using the orbit-averaged approximation (indicated with OA), or locally. In all cases, $M_{\bullet}=3\times 10^{5}\mathrm{M}_{\sun}$ and the MBH is surrounded by a distribution of $m_{\star}=1\mathrm{M}_{\sun}$ stars only.
Despite our attempt to reproduce the system employed in QS23 as close as possible, we cannot replicate their results. Although the shape of the $S(a_{0})$ function is very similar, our $a_{\mathrm{c}}$ is approximately a factor of two smaller than what they found. It is likely that this discrepancy stems from the neglect of the stellar potential in QS23, or in differences in the specific implementation of orbit-averaging. Further investigation is still needed to get a definitive answer. Moreover, we observe that $S(a_{0})$ is not modified by using the 2.5PN term to account for GW radiation instead of the equations detailed in Peters (1964). As discussed in the previous section, for this central MBH mass, the local treatment of two-body relaxation does not have a strong impact on the $S(a_{0})$ function. In fact, we find that in Newtonian dynamics the local treatment results in no discernible deviation of $S(a_{0})$ compared to the orbit-averaged approximation. At second PN order instead, we observe some more significant deviations. In particular, as noted in the previous section, our local approach predicts only slightly more DPs for $a_{0}\simeq 3\times 10^{-3}R_{\mathrm{inf}}$ and more EMRIs for $a_{0}\gtrsim 2\times 10^{-2}R_{\mathrm{inf}}$. Finally, a large difference comes from the choice of dynamics. When going from Newtonian to PN dynamics, $S(a_{0})$ is shifted to larger semi-major axes and cliffhanger EMRIs become more likely. This is because PN orbits can get closer to the MBH without being accreted as DPs ($R_{\rm lc}=5.6R_{\mathrm{g}}$ instead of $8R_{\mathrm{g}}$), thus having the chance of experiencing much larger jumps in the $(a,1-e)$ plane along the loss cone line. This, in turn, means that cliffhanger EMRIs can form from larger $a_{0}$, consistent with Fig. 13.
@@ 4.4 Effect on EMRI rates
[…]
$3\times 10^{5}$ $25$ $3.5\times 10^{-7}$ $4.9\times 10^{-8}$ $6.7\times 10^{-8}$ $3.0\times 10^{-6}$ $2.8\times 10^{-7}$
Table 1: Instantaneous EMRI+DP ($\dot{N}$), EMRI ($\dot{N}_{\mathrm{EMRI}}$) and DP ($\dot{N}_{\mathrm{DP}}$) rates produced in the snapshots we investigated. We report their estimates according to our formulation Eq. (62) (no superscript) based on the transfer function $S(a_{0})$, and the classical estimates Eq. (64) (superscript $\mathrm{cl}$) based solely on the critical semi-major axis $a_{\mathrm{c}}=0.01\,R_{\mathrm{inf}}$.
@@
that can be directly computed during the evolution (Cohn & Kulsrud 1978). The total rate of EMRIs and DPs per unit time can be computed as:
@@
Another possibility is $a_{0}=(r_{-}+r_{+})/2$, but it is affected by the stellar potential at smaller $\mathcal{E}$. In fact the two mappings for typical loss cone orbits depend on the total potential evaluated at $r_{\mathrm{c}}=a_{0}$ and $r_{+}\simeq 2\,a_{0}$ respectively. The rates of DPs and EMRIs can be computed as:
@@
When using the classical discriminant between EMRIs and DPs, one sets $S(\mathcal{E})=\Theta(\mathcal{E}-\mathcal{E}_{\mathrm{c}})$, where $\Theta(x)$ is the Heaviside step function, that is zero when its argument is negative and one otherwise, and $\mathcal{E}_{\mathrm{c}}$ is the energy such that $r_{\mathrm{c}}(\mathcal{E}_{\mathrm{c}})=a_{\mathrm{c}}$. Accordingly, the “classical” rate of the two classes of captures, that we denote with the upper index “cl”, is:
@@
[…]
We first specialise to the case $M_{\bullet}=3\times 10^{5}\,{\rm M}_{\odot}$, which is where the bulk of LISA detections are expected, and use the code presented in Broggi et al. (2022) to simulate the relaxation of the nuclear cluster surrounding it. The initial Dehnen profile of stars and stellar-mass BHs has $\gamma_{\star}=\gamma_{\bullet}=1.5$, while the other properties match those of Sect. 3.1. At later times, the system relaxes and two-body scattering in the inner region are dominated by stellar-mass BHs, whose distribution reaches $\gamma_{\bullet}=1.7$, as in Sect. 3.1. The code accounts for a non-evolving stellar potential set by the initial conditions and includes the effects of two-body relaxation, but does not account for GWs emission. Since the GWs driven region of phase space encloses the loss cone for $a<a_{\mathrm{c}}$, we interpret $\mathcal{F}_{\mathrm{lc}}$ for $\mathcal{E}>\mathcal{E}_{\mathrm{c}}$ as the differential rate entering the GW dominated region [^12] Some recent works refer to the line $t_{\mathrm{GW}}\simeq t_{\mathrm{rlx}}$ as the loss cone for $a<a_{\mathrm{c}}$ (Bondani et al. 2024; Kaur et al. 2024; Rom et al. 2024). In this work, with “loss cone curve” we always refer to the $r_{-}=R_{\mathrm{lc}}$ curve.. For this system $R_{\mathrm{inf}}=0.57$ pc and $\mathcal{E}_{\mathrm{c}}=63\,\sigma^{2}_{\mathrm{inf}}$, where $\sigma_{\mathrm{inf}}=\sigma(R_{\mathrm{inf}})$. We build $S(\mathcal{E})$ by interpolating the function shown in Fig. 9 linearly in the $\log$-$\log$ space. We extrapolate as $S(\mathcal{E})=0$ for $\mathcal{E}\to 0$ ($a_{0}\to\infty$), and $S(\mathcal{E})=1$ for $\mathcal{E}\to\infty$ ($a_{0}\to 0$).
@@
The first two points depend mainly on the fact that $S(a_{0})$ takes approximately an order of magnitude in semi-major axis (or, equivalently, in energy) to smoothly decrease from 1 to its minimum value $S_{\mathrm{min}}$, with $S_{\mathrm{min}}<S(a_{\mathrm{c}})<1$ at any value of $M_{\bullet}$. This shows the limitations of approximating $S(\mathcal{E})$ as a step function. The last point, on the other hand, is exclusively caused by the novelty that $S(a_{0})>0$ even for $a_{0}\gg a_{\mathrm{c}}$ introduced by QS23.
4.4.2 Effects on EMRIs from lower central MBH masses.

### v2

@@ 4.1 EMRI-to-DP ratio and cliffhanger EMRIs
Consider now $a_{\alpha}$ and $a_{\beta}$ as the values of the semi-major axis of the orbit at two subsequent generic apocentres. For each run that resulted in the formation of an EMRI, we consider the largest ratio $a_{\alpha}/a_{\beta}$ to set a criterion to distinguish cliffhanger and traditional EMRIs. Consistently with the reasoning that led us to Eq. (35), we determine the passage across the cliffhanger region if at some point in the orbital evolution the semi-major axis is cut in half or more during a single period, thus for all the runs where $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}>2$. This threshold is somehow arbitrary, but offers a quantitative way to identify cliffhanger EMRIs. Figure 10 shows the result of this investigation for $M_{\bullet}=10^{5}\,\mathrm{M_{\sun}}$ and $M_{\bullet}=10^{6}\,\mathrm{M_{\sun}}$. For both MBH masses, all EMRIs forming for $a_{0}\ll 10^{-2}R_{\mathrm{inf}}$ are classical EMRIs. Both cases show an increase in the average value of $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}$ for larger $a_{0}$, but only for $M_{\bullet}=10^{5}\,\mathrm{M_{\sun}}$ the threshold of 2 is crossed. Moreover, we verify that the prediction for the largest possible $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}$ ratio (which occurs at the loss cone edge) given in Eq. (38) is verified for $M_{\bullet}=10^{5}\,\mathrm{M_{\sun}}$ for large $a_{0}$, with some deviations for smaller $a_{0}$. This is likely because the approximation given in Eq. (37) is less precise for smaller semi-major axes.
@@ 4.2 Local versus orbit-averaged approach
The relaxation timescale for a particle with energy $\mathcal{E}$ and pericentre $r_{-}$ estimates the average time that the particle passes on the orbit before being scattered away. When $r_{-}$ is much smaller than the semi-major axis of the orbit, it estimates the time before $r_{-}$ is changed by the order of itself, as energy remains almost constant (Cohn & Kulsrud 1978). In this small $r_{-}$ regime the relaxation timescale for the orbit can be estimated as $t_{\mathrm{rlx}}\simeq P\,r_{-}/(q\,r_{\mathrm{lc}})$ (Merritt 2013a; Bortolas et al. 2023; Broggi et al. 2024), where $P$ is the orbital period. Consider now an orbit with energy $\mathcal{E}$ and $r_{-}<r_{\mathrm{lc}}$: in a time $P$ it will be reached by a number $\mathrm{d}N$ of stellar objects due to two-body relaxation. In the orbit-averaged approach, they will all be DPs produced in the subsequent orbital period, with a rate from that orbit of $\mathrm{d}\dot{N}=\mathrm{d}N/P$. In the local approach, if $t_{\mathrm{rlx}}<P$, only the stellar objects that will take less than $t_{\mathrm{rlx}}$ to reach the pericentre will be DPs, and the corresponding rate will be $\mathrm{d}\dot{N}\simeq\mathrm{d}N\,t_{\mathrm{rlx}}/P^{2}=r_{-}/(q\,r_{\mathrm{lc}})\,\mathrm{d}N/P$. Therefore, the orbit-averaged approach will overestimate the rate of DPs for $q\gtrsim 1$ due to the ‘frozen orbit’ approach, as they have $r_{-}<r_{\mathrm{lc}}$ by definition. The same reasoning applies to cliffhangers, that are produced at very small pericentres, too. However, being their pericentre larger than $r_{\mathrm{lc}}$ by definition, the overestimation of cliffhangers in the orbit-averaged assumption will be lower. Consequently, we expect that the orbit-averaged approach will underestimate $S(a)$.
@@
In order to test this expectation we re-ran the three sets of simulations with $M_{\bullet}=10^{4},10^{5},3\times 10^{5}\,\mathrm{M}_{\sun}$ with the exact same system parameters but employing the orbit-averaged procedure described in Sec. 3.3. Results of these test runs are compared to our default simulations in Fig. 12, and appear to confirm our expectation. The $S(a_{0})$ functions computed in the local vs orbit-averaged simulations nicely match up to the value of $a_{0}/R_{\rm inf}$ for which $q=1$, which marks the transition from empty to full loss cone (vertical dotted lines in the figure). At larger $a_{0}$, the orbit-averaged approach tends to produce less cliffhanger EMRIs and more plunges, thus resulting in a lower $S(a_{0})$. We note that the trend becomes more severe by lowering the central MBH mass, consistent with the fact that the cliffhanger region becomes larger and larger. In practice, the smaller the central MBH, the bigger gets the cliffhanger region compared to the loss cone, and it becomes increasingly more difficult for the orbit-averaged approximation to capture the detailed dynamics of particle scattered within the cliffhanger region in real time along the orbit close to the pericentre, before getting to the loss cone, thus leading to an over-estimation of DPs.
@@
Figure 12: Comparison of the $S(a_{0})$ function obtained by employing the local (solid lines) vs orbit-averaged (dashed lines) treatment of two-body relaxation. The vertical dotted lines mark, for each case, the value of $a_{0}/R_{\rm inf}$ at which $q=1$, thus marking the transition from the empty (at smaller $a_{0}$) to the full loss cone regime. The crosses indicate the points for which the orbit-averaged procedure explained in Sect. 3.3 failed many times.
4.3 Detailed investigation of the differences with QS24
@@
Our main results display some deviations with respect to other works in the literature. In particular the emergence of cliffhanger EMRIs and their abundance does not match findings by QS24. This is not surprising, given the many differences between our simulations and theirs. In particular, we simulate a two-population system, including the effect of solar-mass stars and stellar-mass BHs in the computation of the diffusion coefficient, we adopt a local treatment of two-body relaxation and we evolve the stellar-mass BH using PN equations of motion. Conversely, in QS24 scattering comes from a single population of solar-mass object, is treated as orbit-averaged and the dynamics of the stellar-mass BH is Newtonian.
@@
In the previous section, we already highlighted the importance of the local treatment of relaxation, which leads to a much larger relative fraction of EMRIs in the full loss cone regime. This can certainly account for some of the differences with QS24. For example in the case of $M_{\bullet}=10^{4}\,\mathrm{M}_{\sun}$, the orbit-averaged tests shown in Fig. 12 result in $S(a_{0})\approx 0.4$ at large $a_{0}$, which is much closer to the value of $\approx 0.25$ found in QS24 compared to the $\gtrsim 0.6$ found in our default simulations employing a local treatment of relaxation.
@@
In order to further test how the differences between the QS24 and our studies affect the results, we perform further targeted simulations, focusing on the $M_{\bullet}=3\times 10^{5}\,\mathrm{M}_{\sun}$ case, as it is close to the MBH mass value at which the majority of LISA EMRI detections is expected (Babak et al. 2017). We try to reproduce the results by QS24111111QS24 do not present results for a $M_{\bullet}=3\times 10^{5}\,\mathrm{M}_{\sun}$ scenario; however, the authors kindly agreed to share the results with us and to let us show them in Fig. 13. by running our simulations with a setup which is as close as possible to theirs. We use a single population of scattering objects of $m_{\star}=1\,\mathrm{M}_{\sun}$, distributed according to a Denhen density profile characterised by $M_{\mathrm{tot},\star}=1425M_{\bullet}$, $r_{\mathrm{a},\star}=223.37R_{\mathrm{inf}}$ and $\gamma_{\star}=1.75$. With these parameters, our density profile departs from the single $r^{-1.75}$ power law simulated by QS24 by less than 1% within the MBH influence radius.
@@
Qunbar & Stone (2024) use Newtonian dynamics, and take into account emission of GWs through the equations detailed in Peters (1964). We do the same by setting $H_{1}=H_{2}=H_{2.5}=0$ and adding the contributions
@@
We report the results of these tests in Fig. 13.
Figure 13: Function $S(a_{0})$ given different assumptions on the dynamics and treatment of two-body relaxation. For the dynamics, we distinguish between Newtonian (N) or PN evolution of the system, and between employing the equations detailed in Peters (1964) or the 2.5PN term to treat loss of energy and angular momentum due to GW emission. Instead, two-body relaxation can be accounted for using the orbit-averaged approximation (indicated with OA), or locally. In all cases $M_{\bullet}=3\times 10^{5}\,\mathrm{M}_{\sun}$ and the MBH is surrounded by a distribution of $m_{\star}=1\,\mathrm{M}_{\sun}$ stars only.
Despite our attempt to reproduce the system employed in QS24 as close as possible, we cannot replicate their results. Although the shape of the $S(a_{0})$ function is very similar, our $a_{\mathrm{c}}$ is approximately a factor of two smaller than what they found. It is likely that this discrepancy stems from the neglect of the stellar potential in QS24, or in differences in the specific implementation of orbit-averaging. Further investigation is still needed to get a definitive answer. Moreover, we observe that $S(a_{0})$ is not modified by using the 2.5PN term to account for GW radiation instead of the equations detailed in Peters (1964). As discussed in the previous section, for this central MBH mass, the local treatment of two-body relaxation does not have a strong impact on the $S(a_{0})$ function. We find that in Newtonian dynamics the local treatment results in no discernible deviation of $S(a_{0})$ compared to the orbit-averaged approximation. At second PN order instead, we observe some more significant deviations. In particular, as noted in the previous section, our local approach predicts only slightly more DPs for $a_{0}\simeq 3\times 10^{-3}R_{\mathrm{inf}}$ and more EMRIs for $a_{0}\gtrsim 2\times 10^{-2}R_{\mathrm{inf}}$. Finally, a large difference comes from the choice of dynamics. When going from Newtonian to PN dynamics, $S(a_{0})$ is shifted to larger semi-major axes and cliffhanger EMRIs become more likely. This is because PN orbits can get closer to the MBH without being accreted as DPs ($R_{\rm lc}=5.6R_{\mathrm{g}}$ instead of $8R_{\mathrm{g}}$), thus having the chance of experiencing much larger jumps in the $(a,1-e)$ plane along the loss cone line. This, in turn, means that cliffhanger EMRIs can form from larger $a_{0}$, consistent with Fig. 13.
@@ 4.4 Effect on EMRI rates
that can be directly computed during the evolution (Cohn & Kulsrud 1978). The total rate of EMRIs and DPs per unit time can be computed as follows:
@@
Another possibility is $a_{0}=(r_{-}+r_{+})/2$, but it is affected by the stellar potential at smaller $\mathcal{E}$. The two mappings for typical loss cone orbits depend on the total potential evaluated at $r_{\mathrm{c}}=a_{0}$ and $r_{+}\simeq 2\,a_{0}$ respectively. The rates of DPs and EMRIs can be computed as
@@
When using the classical discriminant between EMRIs and DPs, one sets $S(\mathcal{E})=\Theta(\mathcal{E}-\mathcal{E}_{\mathrm{c}})$, where $\Theta(x)$ is the Heaviside step function, that is zero when its argument is negative and one otherwise, and $\mathcal{E}_{\mathrm{c}}$ is the energy such that $r_{\mathrm{c}}(\mathcal{E}_{\mathrm{c}})=a_{\mathrm{c}}$. Accordingly, the classical rate of the two classes of captures, which we denote with the upper index ‘cl’, is
@@
We first specialise to the case $M_{\bullet}=3\times 10^{5}\,{\rm M}_{\odot}$, which is where the bulk of LISA detections are expected, and use the code presented in Broggi et al. (2022) to simulate the relaxation of the nuclear cluster surrounding it. The initial Dehnen profile of stars and stellar-mass BHs has $\gamma_{\star}=\gamma_{\bullet}=1.5$, while the other properties match those of Sect. 3.1. At later times, the system relaxes and two-body scattering in the inner region are dominated by stellar-mass BHs, whose distribution reaches $\gamma_{\bullet}=1.7$, as in Sect. 3.1. The code accounts for a non-evolving stellar potential set by the initial conditions and includes the effects of two-body relaxation, but does not account for GWs emission. Since the GWs driven region of phase space encloses the loss cone for $a<a_{\mathrm{c}}$, we interpret $\mathcal{F}_{\mathrm{lc}}$ for $\mathcal{E}>\mathcal{E}_{\mathrm{c}}$ as the differential rate entering the GW dominated region. [^12] Some recent works refer to the line $t_{\mathrm{GW}}\simeq t_{\mathrm{rlx}}$ as the loss cone for $a<a_{\mathrm{c}}$ (Bondani et al. 2024; Kaur et al. 2024; Rom et al. 2024). In this work, with ‘loss cone curve’ we always refer to the $r_{-}=R_{\mathrm{lc}}$ curve. For this system $R_{\mathrm{inf}}=0.57$ pc and $\mathcal{E}_{\mathrm{c}}=63\,\sigma^{2}_{\mathrm{inf}}$, where $\sigma_{\mathrm{inf}}=\sigma(R_{\mathrm{inf}})$. We build $S(\mathcal{E})$ by interpolating the function shown in Fig. 9 linearly in the $\log$-$\log$ space. We extrapolate as $S(\mathcal{E})=0$ for $\mathcal{E}\to 0$ ($a_{0}\to\infty$), and $S(\mathcal{E})=1$ for $\mathcal{E}\to\infty$ ($a_{0}\to 0$).
@@
Table 1: Instantaneous EMRI+DP ($\dot{N}$), EMRI ($\dot{N}_{\mathrm{EMRI}}$), and DP ($\dot{N}_{\mathrm{DP}}$) rates produced in the snapshots we investigated.
$3\times 10^{5}$ $25$ $3.5\times 10^{-7}$ $4.9\times 10^{-8}$ $6.7\times 10^{-8}$ $3.0\times 10^{-7}$ $2.8\times 10^{-7}$
 [^13] We report estimates according to our formulation Eq. (62) (no superscript) based on the transfer function $S(a_{0})$, and the classical estimates Eq. (64) (superscript $\mathrm{cl}$) based solely on the critical semi-major axis $a_{\mathrm{c}}=0.01\,R_{\mathrm{inf}}$.
[…]
@@
The first two points depend mainly on the fact that $S(a_{0})$ takes approximately an order of magnitude in semi-major axis (or, equivalently, in energy) to smoothly decrease from 1 to its minimum value $S_{\mathrm{min}}$, with $S_{\mathrm{min}}<S(a_{\mathrm{c}})<1$ at any value of $M_{\bullet}$. This shows the limitations of approximating $S(\mathcal{E})$ as a step function. The last point, on the other hand, is exclusively caused by the novelty that $S(a_{0})>0$ even for $a_{0}\gg a_{\mathrm{c}}$ introduced by QS24.
4.4.2 Effects on EMRIs from lower central MBH masses